OpenStudy (anonymous):

What is the distance from (-3, 1) to (-1, 5) Round your answer to the nearest hundredth.

OpenStudy (anonymous):

\[distance=\sqrt{(x2-x1)^2+(y2-y1)^2}\] \[\sqrt{(5-1)^2+[-1-(-3)]^2}\] \[\sqrt{(4)^2+(2)^2}\] \[\sqrt{16+4}\] \[\sqrt{20}\] 4.47

OpenStudy (anonymous):

Thanks for explaining it too :)